OpenStudy (anonymous):

Solve for x and check. Enter only a number. 32x = 27

OpenStudy (anonymous):

\[32x=27\] Divide both sides by 32; \[\frac{ 32x }{ 32 }=\frac{ 27 }{ 32 }\] Find out what the value of\[27\div32\]is and that is the value of x.

OpenStudy (anonymous):

And to check; Replace the x with the value you found and when you're done simplifying both sides you should end up with \[27=27\]

OpenStudy (fools101):

32x = 27 X= 27/32 here try this web site you should help you! http://www.mathway.com Let me know if this help you! ~Jess~